Ex 6.3 Class 6 Maths Question 1.
Find:
(a) 35 – (20)
(b) 72 – (90)
(c) (-15) – (-18)
(d) (- 20) – (13)
(e) 23 – (-12)
(f) (-32) – (-40)

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) 35 – (20) = 15 + (20) – (20) = 15 + 0 = 15 [(+a) + (-a) = 0] (b) 72 – 90 72 – (72 + 18) = 72 – 72 – 18 = 0 – 18 = – 18 [a + (- a) = 0] (c) (- 15) – (- 18) = (- 15) + (additive inverse of – 18) = (-15) + (18) = 3 (d) (- 20) – (13) (- 20) – (13) = – [20 + 13] = – 33 (e) 23 – (- 12) 23 – (- 12) = 23 + (additive inverse of – 12) = 23 + 12 = 35 (f) (- 32) – (- 40) (- 32) + (additive inverse of – 40) = (- 32) + 40 = 8